# Hyperbolic Fuzzy $C$-Means with Adaptive Weight-based Filtering for Clustering in Non-Euclidean Spaces

### v1

@@ 6.2 Centroid Update
Table II: Comparison of Clustering Performance across multiple methods, $K$-means, FCM, PCM, P_SFCM, MiniBatchKmeans, HSFC, with our proposed HypeFCM ON $7$ Real-World datasets, presented as Means with Standard Deviations.
Datasets Metric FCM PCM P_SFCM $K$-means MiniBatchKmeans HSFC HypeFCM (Ours)
Iris ARI 0.810 $\pm$ 0.012 0.468 $\pm$ 0.071 0.680 $\pm$ 0.058 0.742 $\pm$ 0.023 0.800 $\pm$ 0.015 0.742 $\pm$ 0.021 0.812 $\pm$ 0.015
NMI 0.815 $\pm$ 0.015 0.548 $\pm$ 0.065 0.725 $\pm$ 0.051 0.756 $\pm$ 0.027 0.788 $\pm$ 0.019 0.756 $\pm$ 0.025 0.817 $\pm$ 0.014
Glass ARI 0.228 $\pm$ 0.045 0.005 $\pm$ 0.010 0.129 $\pm$ 0.022 0.270 $\pm$ 0.041 0.210 $\pm$ 0.047 0.264 $\pm$ 0.052 0.598 $\pm$ 0.007
NMI 0.334 $\pm$ 0.038 0.035 $\pm$ 0.016 0.156 $\pm$ 0.026 0.428 $\pm$ 0.036 0.356 $\pm$ 0.042 0.398 $\pm$ 0.047 0.609 $\pm$ 0.016
Ecoli ARI 0.378 $\pm$ 0.041 0.390 $\pm$ 0.047 0.293 $\pm$ 0.028 0.384 $\pm$ 0.039 0.390 $\pm$ 0.038 0.424 $\pm$ 0.057 0.538 $\pm$ 0.006
NMI 0.472 $\pm$ 0.034 0.349 $\pm$ 0.061 0.324 $\pm$ 0.034 0.534 $\pm$ 0.035 0.545 $\pm$ 0.031 0.547 $\pm$ 0.049 0.591 $\pm$ 0.010
Wine ARI 0.366 $\pm$ 0.032 0.042 $\pm$ 0.019 0.287 $\pm$ 0.015 0.352 $\pm$ 0.042 0.365 $\pm$ 0.036 0.375 $\pm$ 0.038 0.385 $\pm$ 0.006
NMI 0.425 $\pm$ 0.036 0.089 $\pm$ 0.024 0.320 $\pm$ 0.019 0.423 $\pm$ 0.037 0.430 $\pm$ 0.032 0.428 $\pm$ 0.034 0.486 $\pm$ 0.006
Zoo ARI 0.527 $\pm$ 0.025 0.302 $\pm$ 0.054 0.157 $\pm$ 0.067 0.709 $\pm$ 0.022 0.680 $\pm$ 0.044 0.447 $\pm$ 0.062 0.726 $\pm$ 0.003
NMI 0.543 $\pm$ 0.029 0.457 $\pm$ 0.049 0.376 $\pm$ 0.058 0.782 $\pm$ 0.025 0.747 $\pm$ 0.039 0.482 $\pm$ 0.058 0.788 $\pm$ 0.005
Abalone ARI 0.144 $\pm$ 0.062 $0$$\pm$ 0.002 0.161 $\pm$ 0.019 0.106 $\pm$ 0.069 0.129 $\pm$ 0.059 0.103 $\pm$ 0.033 0.162 $\pm$ 0.010
NMI 0.134 $\pm$ 0.057 0.002 $\pm$ 0.002 0.163 $\pm$ 0.027 0.099 $\pm$ 0.072 0.126 $\pm$ 0.053 0.097 $\pm$ 0.038 0.159 $\pm$ 0.009
Phishing (5k) ARI $0$$\pm$ 0.003 $0$$\pm$ 0.015 0.060 $\pm$ 0.046 0.001 $\pm$ 0.001 0.001 $\pm$ 0.002 0.002 $\pm$ 0.001 0.169 $\pm$ 0.003
NMI 0.010 $\pm$ 0.005 0.003 $\pm$ 0.008 0.069 $\pm$ 0.031 0.001 $\pm$ 0.007 0.003 $\pm$ 0.001 0.005 $\pm$ 0.002 0.207 $\pm$ 0.002
    Average Rank 3.86 6.29 4.54 4.09 3.30 3.64 1.12
@@
7.6 Ablation Studies.
We conducted an ablation study by varying the key parameter Filtration.
Figure 5: Comparison of Clustering Performance, NMI vs $k$ with and without filtration for two real-world (a) Wine, (b) Ecoli and two synthetic (c) cure-t1-2000n-2D, (d) smile1 datasets respectively.
HypeFCM with and without Filtration. We conducted two separate experiments, one with filtration and the other without using any filtration on two real-world datasets, Wine, Ecoli and two synthetic datasets, cure-t1-2000n and smile1 respectively. We have performed these experiments by setting the curvature parameter $(\alpha)$ as $1$. Figure 5 shows that using filtration improves the performance of HypeFCM compared to the same without filtration.
Table III: Details of the Datasets Used for the Experimentation.
@@
8 Discussions.
The performance results of HypeFCM demonstrate its effectiveness across both synthetic and real-world datasets. The algorithm’s superior performance in terms of ARI and NMI metrics against competing methods validates the fundamental hypothesis that hyperbolic geometry can better capture complex hierarchical relationships in clustering tasks.
The significant performance of HypeFCM on $7$ real-world benchmark datasets, as visualized through t-SNE for Glass, Ecoli and Wine datasets, demonstrates how effectively HypeFCM preserves cluster boundaries and captures the inherent hierarchical structures that often characterize real-world datasets. Integrating hyperbolic geometry via the Poincaré Disc model represents a significant methodological advancement. The model’s exponentially expanding space toward its boundary naturally accommodates hierarchical structures that are prevalent in many real-world datasets but challenging to represent in Euclidean space. This geometric advantage, combined with fuzzy clustering principles, allows HypeFCM to capture subtle hierarchical relationships that traditional fuzzy clustering methods cannot properly identify.
9 Conclusion & Future works.
Our proposed HypeFCM marks a significant advancement in handling complex hierarchical data structures by seamlessly integrating hyperbolic geometry with the principles of fuzzy clustering. Our method exploits the unique properties of the Poincaré Disc model, which provides exponentially more space toward its boundary, making it naturally suited for representing hierarchical relationships. Incorporating a weight-based filtering mechanism enhances the algorithm’s ability to focus on relevant geometric relationships while pruning less significant connections, thereby improving both computational efficiency and clustering accuracy. The algorithm’s iterative optimization process, combining hyperbolic distance calculations with fuzzy membership updates, demonstrates robust performance in capturing complex structures that traditional Euclidean-based methods struggle to represent adequately. Experimental results validate the effectiveness of our approach, showing superior performance in preserving hierarchical relationships while maintaining computational feasibility. While the current results are promising, expanding this framework to handle streaming data and exploring integration with deep learning architectures may further enhance the algorithm’s applicability across diverse domains where hierarchical relationships are intrinsic to the data structure.

### v2

@@ 7.5 Ablation Studies
Table III: Comparison of Clustering Performance across multiple Fuzzy contenders, FCM, PCM, P_SFCM, EFKM, UFCER, FCSR, IFKMHC, HSFC, FCPFM, with our proposed HypeFCM, presented as Means with Standard Deviations. Best and second-best results are highlighted in bold and underlined, respectively.
Dataset Metric FCM PCM P_SFCM EFKM UFCER FCSR IFKMHC HSFC FCPFM HypeFCM (Ours)
[…]
Zoo ARI 0.527 $\pm$ 0.025 0.302 $\pm$ 0.054 0.157 $\pm$ 0.067 0.343 $\pm$ 0.015 0.675 $\pm$ 0.015 0.591 $\pm$ 0.015 0.611 $\pm$ 0.015 0.447 $\pm$ 0.062 0.671 $\pm$ 0.035 0.726 $\pm$ 0.018
NMI 0.543 $\pm$ 0.029 0.457 $\pm$ 0.049 0.376 $\pm$ 0.058 0.427 $\pm$ 0.015 0.711 $\pm$ 0.015 0.615 $\pm$ 0.015 0.664 $\pm$ 0.015 0.482 $\pm$ 0.058 0.723 $\pm$ 0.045 0.788 $\pm$ 0.025
[…]
Average Rank $6.967$ $9.950$ $6.783$ $4.900$ $6.633$ $3.300$ $6.783$ $3.800$ $4.133$ $\mathbf{1.466}$
@@
Lemma 4.
For any two points $\mathbf{x},\mathbf{y}\in\mathbb{D}_{\alpha}^{p}$, the function $d_{hyp}(\mathbf{x},\mathbf{y})$ is geodesically convex function in $\mathbb{D}_{\alpha}^{p}$.
Proof.
Let us prove that the hyperbolic distance function $d_{hyp}$ is geodesically convex by showing that for any geodesic $\bm{\gamma}(t)$ in $\mathbb{D}^{p}_{\alpha}$, the function $t\mapsto d_{hyp}(\mathbf{x},\bm{\gamma}(t))$ is convex for any fixed point $\mathbf{x}$.
Let $\bm{\gamma}:[0,1]\to\mathbb{D}^{p}_{\alpha}$ be a geodesic. Define:
$f(t)=d_{hyp}(x,\bm{\gamma}(t))=\frac{2}{\sqrt{\alpha}}\tanh^{-1}(\sqrt{\alpha}\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|).$ (26)
Computing the first derivative using the chain rule:
$f^{\prime}(t)$ $=\frac{2}{\sqrt{\alpha}}\frac{1}{1-\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|^{2}}\frac{d}{dt}\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|.$ (27)
The second derivative is:
$f^{\prime\prime}(t)$ $=\frac{2}{\sqrt{\alpha}}\Bigg{[}\frac{2\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|}{(1-\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|^{2})^{2}}\Big{(}\frac{d}{dt}\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|\Big{)}^{2}$ (28)
$\quad+\frac{1}{1-\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|^{2}}\frac{d^{2}}{dt^{2}}\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|\Bigg{]}.$
Now we have :
1. $\frac{2\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|}{(1-\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|^{2})^{2}}$ is positive because:
  • $\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|>0$ for $\mathbf{x}\neq\bm{\gamma}(t),$
  • $1-\alpha\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|^{2}>0$ due to the properties of the Poincaré disc model.
2. $\Big{(}\frac{d}{dt}\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|\Big{)}^{2}$ is non-negative.
3. $\frac{d^{2}}{dt^{2}}\|-\mathbf{x}\oplus_{\alpha}\bm{\gamma}(t)\|$ is non-negative along geodesics due to the negative curvature of the hyperbolic space. Since all terms in $f^{\prime\prime}(t)$ are non-negative, we have:
$f^{\prime\prime}(t)\geq 0.$
Therefore, $f(t)$ is convex along any geodesic $\bm{\gamma}(t)$, proving that $d_{hyp}$ is geodesically convex. ∎
Proof of Lemma 2.
Proof.
The objective is geodesically convex by Lemma 4. The domain $\mathbb{D}_{\alpha}^{p}$ is a complete metric space by Lemma 3. By standard optimization theory in metric spaces, specifically the Existence and Uniqueness theorem for convex optimization in complete metric spaces, $f$ attains its minimum on $\mathbb{D}_{\alpha}^{p}$. The minimizer is unique due to strict geodesic convexity. We denote this unique minimizer as $\mathbf{v}_{j}^{*}$. ∎
Table IV: Statistical Significance of Clustering Methods Compared to HypeFCM (ours) via $p$-values for ARI and NMI Metrics.
Datasets Metric FCM PCM P_SFCM $K$-means MiniBatchKMeans IFKMHC EFKM UFCER FCSR HSFC FCPFM
Cure-t1-2000n ARI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 2.2811e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 8.6000e-01 6.1000e-05 4.1210e-02 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Cure-t2-4k ARI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 7.1700e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Donutcurves ARI 6.1000e-05 7.2200e-04 8.1000e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 7.1450e-02 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 4.3100e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-01 6.1000e-05 6.1000e-05
Smile1 ARI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 3.3100e-01 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.3250e-02 6.1000e-05 6.1000e-05
Disk-4000n ARI 2.0500e-03 5.6000e-05 1.0800e-04 5.6100e-03 5.6100e-03 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 4.6000e-03 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
3MC ARI 1.0800e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 1.1100e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Iris ARI 7.1320e-02 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 8.1571e-02 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Glass ARI 6.1000e-05 2.9000e-02 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Ecoli ARI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Wine ARI 3.1000e-04 6.1000e-05 2.9000e-05 6.1000e-05 8.1000e-04 6.1000e-05 7.1000e-02 6.1000e-05 6.1000e-05 6.1000e-05 1.1110e-01
NMI 4.0100e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 1.1000e-01 6.1000e-05 6.1000e-05 2.1000e-02 5.1020e-02
Zoo ARI 6.1000e-05 6.1000e-05 6.1000e-05 1.1120e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 2.1000e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Abalone ARI 7.1500e-04 6.1000e-05 1.0800e-01 8.8300e-04 7.1500e-04 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 8.2100e-02
NMI 6.1000e-05 6.1000e-05 1.1000e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1022e-02
Phishing (5k) ARI 6.1000e-05 6.1000e-05 6.1000e-05 8.2100e-03 2.7400e-03 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 1.1800e-03 1.1000e-01
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 8.1201e-02
Wisconsin B.C. ARI 2.9000e-03 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.5000e-02 5.1230e-01
NMI 7.1000e-03 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 7.1000e-01 3.1025e-01
Flights (2k) ARI 1.1900e-03 7.1000e-04 2.9000e-03 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
Flights (5k) ARI 6.1000e-05 2.0500e-03 8.2300e-03 6.1000e-05 6.1000e-05 6.1000e-05 8.8100e-02 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 2.1100e-02 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05
MNIST (5k) ARI 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 2.1000e-01 6.1000e-05 6.1000e-05
NMI 6.1000e-05 6.1000e-05 1.0700e-01 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 4.1000e-01 6.1000e-05 6.1000e-05
ORHD ARI 6.1000e-05 7.1900e-04 2.2200e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 2.1020e-01 1.0800e-01 6.1000e-05
NMI 6.1000e-05 6.1000e-05 7.1000e-04 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1000e-05 6.1560e-02 1.1000e-01 6.1000e-05
Figure 7: t-SNE plots on Ecoli datasets for different Iterations for FCM and HypeFCM methods.
Appendix B More Experimentation on Datasets.
We have compared our proposed method, HypeFCM, with other Fuzzy contenders in our paper, FCM, PCM, P-SFCM, EFKM, UFCER, FCSR, IFKMHC, HSFC, and FCPFM, on additional datasets.
We have noted the performances of the datasets in Table III. The results in Table III indicate that our proposed method, HypeFCM, outperforms almost all the fuzzy benchmark contenders across all datasets in terms of ARI [45] and NMI [46] values. The t-SNE [50] visualization of HypeFCM in Figure 6 of $5$ real-world datasets, Flights (5k), Wisconsin Breast Cancer, ORHD, Zoo, Phishing URL (5k), demonstrates how effectively HypeFCM preserves cluster boundaries and captures the inherent hierarchical structures of these datasets. The Average Rank in Table III is just the average of the ranks of ARI, and NMI values obtained from the executions of each method $15$ times, all initialized with the same seeding at a time.
In Table IV, we have shown the statistical significance of the methods compared to our proposed HypeFCM for each dataset via $p$-values for the metrics, ARI and NMI. The Wilcoxon Signed Rank [51] statistic is computed by ranking the absolute differences between paired observations of the ARI and NMI values individually for each of the datasets, by contrasting each method $15$ times, all initialized with the same seeding. The final test statistic is the smaller of these sums, and the $p$-value is derived from its distribution to assess whether the method differs significantly from HypeFCM. The lower $p$-values ($p\leq 0.05$) indicate that a statistically significant difference in the performance of the clustering method with our proposed HypeFCM, denoted by or else denotes that there is no significant difference.
We have also tested our method with the other contenders for image segmentation. For the HypeFCM algorithm, we select the fuzzy weighting exponent $m$ = 2, and curvature parameter $\alpha$ = 1. Three images are taken from the BSDS500 dataset, #3096,#12003,#134052,\#3096,\#12003,\#134052, and the Cameraman image dataset from https://github.com/antimatter15/cameraman are used for image segmentation, shown in figure 5.
Table V: Details of all Datasets used for the Experimentation.
@@
Wisconsin Breast Cancer 699 9 2
Abalone 4177 8 2
ORHD 5620 64 10
MNIST 60000 784 10